# Reinforcement Learning for Self-Improving Agent with Skill Library

### v1

@@ Appendix F Expert Experience Dataset Generation
Using these collected examples as an expert experience dataset, we employed Llama-Factory (zheng2024llamafactory) to perform full parameter fine-tuning with 4xNVIDIA H100 8-GPU nodes. During fine-tuning, since all trajectories in the expert experience dataset consist of multiple turn interactions, we did gradient update exclusively on the agent responses with prompt and environment outputs masked. The model was trained with a batch size of 128 and employed a learning rate of 1e-6, using a cosine scheduling strategy with 0.1 warmup ratio. Due to Llama-Factory’s limitations in sequence parallel, we set the maximum token length to 12,500, and any examples exceeding this limit were excluded from the dataset.
@@ I.1 Details of Retrieval Methods
Query N-gram. The core concept of this retrieval method is to leverage skills generated from tasks with similar queries xu2025survey. Under this approach, each query is associated with its own skill library, which can be directly inherited once being retrieved. For simplicity, we initially implement a model-free retrieval approach based on 2-gram Jaccard Similarity to identify the most similar query. To ensure retrieval quality, we establish a threshold of 0.5 for the Jaccard Similarity score; any retrieved queries falling below this threshold are discarded.
Query Embedding. This retrieval method shares the same fundamental principle as Query N-gram, but employs a more advanced text-embedding model to compute embedding similarity for query retrieval. Specifically, we utilize the all-MiniLM-L6-v2 model (reimers-2019-sentence-bert) for computing query embeddings with a threshold of 0.65 for the cosine similarity.
Skill Embedding. Skill Embedding retrieves relevant skills from the skill library for each task query using a standard retrieval model. Each newly generated skill is encoded into embeddings and indexed in the skill library for future retrieval purposes. We employ Qwen3-Embedding-0.6B (qwen3embedding) as our retrieval model and keep the top 5 retrieved skills for usage. This model differs from the general text-embedding model used for Query Embedding because it is specifically trained for document retrieval, where we treat skills as documents and task instructions as queries.

### v2

@@ Appendix F Expert Experience Dataset Generation
Using these collected examples as an expert experience dataset, we employed Llama-Factory (Zheng et al., 2024c) to perform full parameter fine-tuning with 4xNVIDIA H100 8-GPU nodes. During fine-tuning, since all trajectories in the expert experience dataset consist of multiple turn interactions, we did gradient update exclusively on the agent responses with prompt and environment outputs masked. The model was trained with a batch size of 128 and employed a learning rate of 1e-6, using a cosine scheduling strategy with 0.1 warmup ratio. Due to Llama-Factory’s limitations in sequence parallel, we set the maximum token length to 12,500, and any examples exceeding this limit were excluded from the dataset.
@@ I.1 Details of Retrieval Methods
Query N-gram. The core concept of this retrieval method is to leverage skills generated from tasks with similar queries Xu et al. (2025). Under this approach, each query is associated with its own skill library, which can be directly inherited once being retrieved. For simplicity, we initially implement a model-free retrieval approach based on 2-gram Jaccard Similarity to identify the most similar query. To ensure retrieval quality, we establish a threshold of 0.5 for the Jaccard Similarity score; any retrieved queries falling below this threshold are discarded.
Query Embedding. This retrieval method shares the same fundamental principle as Query N-gram, but employs a more advanced text-embedding model to compute embedding similarity for query retrieval. Specifically, we utilize the all-MiniLM-L6-v2 model (Reimers and Gurevych, 2019) for computing query embeddings with a threshold of 0.65 for the cosine similarity.
Skill Embedding. Skill Embedding retrieves relevant skills from the skill library for each task query using a standard retrieval model. Each newly generated skill is encoded into embeddings and indexed in the skill library for future retrieval purposes. We employ Qwen3-Embedding-0.6B (Zhang et al., 2025) as our retrieval model and keep the top 5 retrieved skills for usage. This model differs from the general text-embedding model used for Query Embedding because it is specifically trained for document retrieval, where we treat skills as documents and task instructions as queries.